t(min) [A](M)
21.0 1.260
124.0 9.700X10-1
385.0 4.995X10-1
673.0 2.401X10-1
922.0 1.275X10-1
a.) What is the order of the reaction with the above data?

b.) Calculate the k for the reaction.

c.) Calculate the original concentration of A for the reaction.

d.) Assuming the original concentration in part (c) calculate the half-life of the reaction.
1
Expert's answer
2020-04-15T01:52:34-0400

Solution.

1.

This figure shows graphs of concentration-time dependencies for 0, 1, 2, and 3 reaction orders. As we can see, a linear dependence of order 1 is observed for these values.



2.

"k = \frac{1}{t} \times ln(\frac{C0}{C})"

k = 0.0025

3.

"C0 = e^{kt} \times C"

"C0 = e^{0.0025 \times 21} \times 1.26 = 1.33 \ M"

4.

"t(1/2) = \frac{ln(2)}{k}"

t(1/2) = 277.26 min

Answer:

1.

1 order

2.

k = 0.0025

3.

1.33 M

4.

t(1/2) = 277.26 min